Monochromatic light falling on two slits 0.016 mm
Answer #1

the nth fringe can be found by

n*lambda/d = x/L

where
n = the fringe order (=5)
lambda = wavelength
d = separation of slits (=1.6e-2 mm)
x = distance to the fringe from the zero order
L = distance from the slits to the screen

since you know the angle of the position of the 5th order slit, you can determine x/L by tan(theta)= x/L

Solve

5 * lambda / (1.6e-2) = tan(8.8 deg)

lamda = 4.953 mm...........Ans.